# A SAM-guided Two-stream Lightweight Model for Anomaly Detection

### v1

@@ 4.2.2. Evaluations on other more difficult benchmarks
To conduct more comprehensive evaluations of the anomaly detection capabilities, we further subject our networks to benchmarking using two additional widely recognized datasets VisA (Zou et al., 2022) and DAGM (Zavrtanik et al., 2021a). To provide a comprehensive overview of anomaly localization capabilities, we include 4 metrics: image-level AUROC, pixel-level AUROC, PRO and AP. The results, as presented in Table 2, indicate that our method achieves SOTA pixel-level PRO and AP metrics of 93.83/47.63 on the VisA dataset. Our method has also demonstrated its superiority on the DAGM dataset. On average, STLM performs exceptionally well, affirming its effectiveness and scalability when faced with increasingly complex datasets.
@@ 4.3.1. Visualization on MVTec
To further analyze the model performance, we visualize a qualitative evaluation of anomaly localization performance presented in Figure 7. The visual results demonstrate our model’s precise localization of anomalies. However, no one is perfect. In Figure 9, we analyze instances where our method does not perform as expected. As evidenced in the first and second columns, the susceptibility of our FA module holds the primary responsibility. But as mentioned in (Zhang et al., 2023c), we must acknowledge that several uncertain ground truths are accountable for the observed failures. Considering the results in the third and fourth columns where the ground truth highlights the designated location, our masks only cover the anomalous region on the same images.
@@ 4.4.1. The importance of Large Teacher, Plain Student Stream, Mask Decoder, and Feature Aggregation
In Table 4, we assess the effectiveness of our four design components by conducting experiments where we eliminate the fixed teacher network and use a pre-trained mobileSAM as the teacher network to examine the contribution of SAM-knowledge, remove the plain student stream and use the fixed SAM as the plain stream in the TLM like early KD frameworks (but still use both normal and pseudo anomaly images to train denoising student stream), employee the image embedding exacted from the encoders as the FA module input and train with the cosine loss between paired outputs of two encoders, and substitute the FA module with an empirical feature fusion strategy (Wang et al., 2021). The best results are achieved when all four key design components are combined.
Notably, our work outperforms the “w/o-PlainS” manner, explaining that the plain student stream can learn generalized knowledge related to anomaly detection tasks. This enables our plain student stream to effectively represent previously unseen patterns, ensuring differences between features of two streams are well-captured.
@@
Large Teacher Plain Student Stream Mask Decoder Feature Aggregation I-AUROC P-AUROC PRO
@@
Table 5. Ablation studies on training strategy between one-stage and two-stage training on the MVTec dataset (Bergmann et al., 2019a).
@@
4.4.2. Effect for One-stage training strategy
We conduct ablation studies to investigate the impact of two distinct training strategies mentioned in Section 3.3, as detailed in Table 5. The results reveal that one-stage training delivered superior performance compared to separate training of the TLM and FA module proposed in (Zhang et al., 2023c), particularly in terms of PRO scores. This suggests that training the TLM with additional supervised signals of the pseudo anomaly mask $M$ enhances the network’s ability to locate tiny defects or accurately delineate defect boundaries.
@@ 4.4.2. Effect for One-stage training strategy
STLM (Ours) 99.05 98.26 94.92
@@ 4.4.3. Discrepancy among different manners for the input of the Feature Aggregation module
Figure 10. Experience on the probability of Pseudo Anomaly Generation being activated.
The first way computes the cosine distance between the feature maps of our two-stream lightweight model, making use of more prior information from STLM, which is trained through optimization of the cosine loss function. The second way involves concatenating features of the normal image and the anomalous residual representation defined by (Zhang et al., 2023d). The anomalous feature residuals denote the element-wise Euclidean distance between tensors from two streams. The third way directly concatenates the feature maps $S_{P}^{k}$ and $S_{D}^{k}$ of TLM as the input of our FA module.
The second and the third ways preserve the information from TLM more effectively. The results are presented in Table 6, indicating our method maintains a balance between prior knowledge and representation of information.
@@ 4.4.4. Effect of the knowledge distillation framework
We investigate the importance of the fixed SAM teacher, i.e., only utilizing TLM during training and inference phases, as illustrated in Figure 12. The results are shown in Table 7. We find that the knowledge distillation framework, comparing the two-stream representations, is indispensable for detecting anomalous regions.
4.4.5. Analysis of the probability of Pseudo Anomaly Generation
The probability of the Pseudo Anomaly Generation process being activated governs the extent to which the training datasets deviate from the original normal datasets. To be specific, a higher probability results in a larger proportion of abnormal images and a smaller proportion of normal images. The probability of 0.5 denotes an optimal balance between pseudo-anomalous images and normal images, as shown in Figure 10, achieving the best performance. Since the “good” samples in the test dataset, the probability of 1.0 lets normal images be “unseen” images during training, leading to a high false negative. We design experiments to verify this explanation in Section 4.4.6
@@ 4.4.5. Analysis of the probability of Pseudo Anomaly Generation
In Table 10, we provide further quantitative results of the false negative rate (FNR) (Altman and Bland, 1994) to verify the analysis in the ablation study that the probability of 1.0 lets normal images be “unseen” images during training, leading to a high false negative. The false negative rate, i.e., miss rate, is defined to measure the proportion of instances that are incorrectly identified as negative when it is actually positive. The FNR is calculated as:
@@
Figure 11. Ablation studies on the feature extracted from the $k$-th layer of mask decoder.
In Equation (3) and Equation (4), we define the layers of the feature extracted from the Mask Decoder as the $k$-th layer, which is also presented in Figure 2: Element-wise Production. We conducted a series of experiments to investigate the impact of $k$, and the corresponding results are presented in Figure 11. We denoted the two layers as L1 and L2. We include image-level AUROC, pixel-level AUROC and PRO metrics to provide a comprehensive cooperation of the anomaly detection and localization capabilities of each layer.
@@ 4.4.7. Ablation studies on the feature extracted from the $k$-th layer of mask decoder
Class w/o fixed SAM STLM(Ours)
I↑ P↑ O↑ I↑ P↑ O↑
Carpet 87.92 91.13 69.53 99.48 99.91 96.97
Grid 98.50 97.10 87.62 95.57 95.40 91.56
Leather 99.93 98.30 98.22 100 99.11 98.05
Tile 100 98.98 98.45 100 99.58 99.23
Wood 95.97 91.24 85.26 100 97.99 98.41
Avg. Text. 96.46 95.35 87.82 99.01 98.40 96.84
Bottle 74.13 76.65 74.63 100 97.77 90.92
Cable 69.11 85.15 58.32 98.98 96.95 92.11
Capsule 77.45 87.00 62.05 98.69 98.48 97.28
Hazelnut 86.00 95.27 86.86 100 99.69 98.29
Metal_nut 98.54 83.56 78.07 100 99.45 92.95
Pill 85.45 86.66 87.84 97.20 98.61 94.69
Screw 99.63 96.18 87.44 99.47 97.56 98.60
Toothbrush 87.39 93.25 65.65 99.06 99.61 96.14
Transistor 93.29 79.09 67.94 97.58 94.81 82.50
Zipper 87.66 90.75 79.95 99.74 98.93 96.15
Avg. Obj. 85.87 87.36 74.88 99.07 98.19 93.96
Average 89.40 90.02 79.19 99.05 98.26 94.92
[…]
Method Avg. Text. Avg. Obj. Average
I↑ P↑ O↑ E↓ I↑ P↑ O↑ E↓ I↑ P↑ O↑ E↓
w/o initial weight 99.76 98.87 97.63 129 98.51 98.18 93.03 106 98.93 98.41 94.57 141
STLM (Ours) 99.01 98.40 96.84 38 99.07 98.19 93.96 63 99.05 98.26 94.92 55
[…]
Method I-AUROC P-AUROC PRO
Logit-based 93.45 91.65 80.38
Feature-based 99.05 98.26 94.92
Method Avg. Text.↓ Avg. Obj.↓ Average↓
@@
Cutpaste(Parch) 84.74 89.85 82.03
Cutpaste(Scar) 96.63 96.30 90.52
Cutpaste(Union) 97.49 97.16 91.06
4.4.8. Effect of the initial weight of TLM’s image encoder
We compare the results of employing pre-trained parameters with MobileSAM (Zhang et al., 2023a) and utilizing randomly initialized parameters for ViT-Tiny (Wu et al., 2022) in our TLM, shown in Table 8. We observe that a pre-trained image encoder allows our method to achieve optimal results in fewer epochs. At first glance, the two methods seem comparable for AD tasks. However, upon closer inspection, the ”w/o initial weight” method outperforms in texture classes, while our STLM is better suited for object classes. This leads us to analyze in future research whether SAM’s prior knowledge, centered around object segmentation, plays a role in its superior performance in object-related tasks, despite its less effective representation in texture classes.
4.4.9. Ablation studies on the distillation methods
SAM generates high-quality segmentation results for previously unseen images, making it suitable for logit distillation. As (Zhao et al., 2022) said, distillation methods often focus on distilling deep features from intermediate layers, while the significance of logit distillation is greatly overlooked. We compare the feature-based method with the logit-based method, which minimizes the focal loss (Lin et al., 2017) and L1 loss (Girshick, 2015) between segmentation masks of teacher SAM and the student streams (13). The experimental results, shown in Table 10, indicate that the performance of feature distillation is superior on average, compared with directly utilizing segmentation results from a SAM-like architecture.
@@ 4.4.9. Ablation studies on the distillation methods
To evaluate the pseudo anomaly generation strategies, we present experimental results with the renowned strategies NSA (Schlüter et al., 2022) and Cutpaste (Li et al., 2021b) in Table 11. In most SOTA methods, i.e., DRAEM (Zavrtanik et al., 2021a), Prototypical (Zhang et al., 2023d) and DeSTSeg (Zhang et al., 2023c), the authors attempt to leverage synthetic anomalies generated by NSA, which are also utilized in our paper. Furthermore, based on NSA and Cutpaste,  (Kim et al., 2024; Yao et al., 2023b; Zhang et al., 2023d; Tien et al., 2023) employ a variety of complex or artificial operations that more closely resemble the real anomalies. Notably, SimpleNet (Liu et al., 2023) simulates anomalies by introducing Gaussian noise to normal features, a process that requires careful tuning of appropriate hyperparameters.
@@ 5. Conclusion
In industrial anomaly detection, the primary concerns in real-world applications include two aspects: model efficiency and mobile-friendliness. In this paper, we propose a novel framework called SAM-guided Two-stream Lightweight Model for unsupervised anomaly detection tasks, tailored to meet the demands of real-world industrial applications while capitalizing on the strong generalization ability of SAM. Our STLM effectively distills distinct knowledge from SAM into the Two-stream Lightweight Model (TLM), assigning different tasks to each stream. To be specific, one stream focuses on generating discriminative and generalized feature representations in both normal and anomalous locations, while the other stream reconstructs features without anomalies. Last, we employ a shared mask decoder and a feature aggregation module to generate anomaly maps. By the design of our Two-stream Lightweight Model, shared mask decoder and lightweight feature aggregation module, the calculation and the number of parameters of the networks can be significantly reduced and the inference speed can naturally be improved without noticeable accuracy loss. Experiments on the renowned dataset MVTec AD and two more difficult benchmarks VisA and DAGM demonstrate that our method achieves competitive results on anomaly detection tasks. Furthermore, it exhibits model efficiency and mobile-friendliness compared to state-of-the-art methods.

### v2

@@ 4.2.2. Evaluations on other more difficult benchmarks
To conduct more comprehensive evaluations of the anomaly detection capabilities, we further subject our networks to benchmarking using three additional widely recognized datasets VisA (Zou et al., 2022), MVTec LOCO (Bergmann et al., 2022) and DAGM (Zavrtanik et al., 2021a). VisA features different objects in more types, incorporating three diverse types: complex structures, multiple instances, and single instances. MVTec LOCO contains images from five different categories inspired by real-world industrial inspection scenarios. Structural anomalies appear as scratches, dents, or contaminations in the manufactured products. DAGM contains 10 textured objects with smaller abnormal regions that bear a strong visual resemblance to the background. To demonstrate the intricacy of these three datasets, we present a comparison with the MVTec dataset in Figure 3. Furthermore, the results across all methods reveal a consistent decline for these three datasets.
To provide a comprehensive overview of anomaly localization capabilities, we include 4 metrics: image-level AUROC, pixel-level AUROC, PRO and AP. The results, as presented in Table 3, indicate that our method achieves SOTA pixel-level PRO and AP metrics of 93.83/47.63 on the VisA dataset and SOTA image-level AUROC and pixel-level PRO metrics of 93.18/90.51 on the VisA dataset. Our method has also demonstrated its superiority on the DAGM dataset. On average, STLM performs exceptionally well, affirming its effectiveness and scalability when faced with increasingly complex datasets.
4.2.3. Evaluations on model parameters and inference flexibility
We have conducted additional experiments regarding model parameters and inference flexibility, comparing our method with those mentioned in our paper. The specific results are presented in  5. Overall, our method achieves superior results compared to other methods on average across four datasets. Additionally, our method demonstrates significant advantages in terms of the number of parameters and GPU memory usage, which is highly valuable for practical industrial applications. Moreover, our method achieves the second-fastest speed and ranks third in GFLOPS, making it the best overall in terms of both performance and efficiency.
@@ 4.3.1. Visualization on MVTec
To further analyze the model performance, we visualize a qualitative evaluation of anomaly localization performance presented in Figure 4. The visual results demonstrate our model’s precise localization of anomalies. However, no one is perfect. In Figure 6, we analyze instances where our method does not perform as expected. As evidenced in the first and second columns, the susceptibility of our FA module holds the primary responsibility. But as mentioned in (Zhang et al., 2023b), we must acknowledge that several uncertain ground truths are accountable for the observed failures. Considering the results in the third and fourth columns where the ground truth highlights the designated location, our masks only cover the anomalous region on the same images.
@@ 4.4.1. The importance of Large Teacher, Plain Student Stream, Mask Decoder, and Feature Aggregation
In Table 6, we assess the effectiveness of our four design components by conducting experiments where we eliminate the fixed teacher network and use a pre-trained mobileSAM as the teacher network to examine the contribution of SAM-knowledge, remove the plain student stream and use the fixed SAM as the plain stream in the TLM like early KD frameworks (but still use both normal and pseudo anomaly images to train denoising student stream), employee the image embedding exacted from the encoders as the FA module input and train with the cosine loss between paired outputs of two encoders, and substitute the FA module with an empirical feature fusion strategy (Wang et al., 2021). The best results are achieved when all key design components are combined.
Notably, our work outperforms the “w/o-PlainS” manner, explaining that the plain student stream can learn generalized knowledge related to anomaly detection tasks. Based on the results from Section 4.4.2, where one-stage training outperforms two-stage training, we believe this is due to the FA and TLM modules being trained simultaneously. The segmentation knowledge related to anomaly detection images is transmitted forward by the loss of the FA module, helping both students better recognize normal regions. Without distillation training for SAM, only the denoising student is improved, resulting in differences in normal regions and affecting the experimental results.This enables our two student streams to effectively represent previously unseen patterns, ensuring differences between features of two streams are well-captured.
@@
Large Teacher Plain Student Mask Decoder FA I-AUROC P-AUROC PRO
@@
Table 7. Ablation studies on training strategy between one-stage and two-stage training.
@@
Ours 99.05 98.26 94.92
4.4.2. Effect for One-stage training strategy
We conduct ablation studies to investigate the impact of two distinct training strategies mentioned in Section 3.3, as detailed in Table LABEL:tab:tab5. The results reveal that one-stage training delivered superior performance compared to separate training of the TLM and FA module proposed in (Zhang et al., 2023b), particularly in terms of PRO scores. This suggests that training the TLM with additional supervised signals of the pseudo anomaly mask $M$ enhances the network’s ability to locate tiny defects or accurately delineate defect boundaries.
@@ 4.4.3. Discrepancy among different manners for the input of the Feature Aggregation module
The first way computes the cosine distance between the feature maps of our TLM, making use of more prior information from STLM, which is trained through optimization of the cosine loss function. The second way involves concatenating features of the normal image and the anomalous residual representation defined by (Zhang et al., 2023c). The anomalous feature residuals denote the element-wise Euclidean distance between tensors from two streams. The third way directly concatenates the feature maps $S_{P}^{k}$ and $S_{D}^{k}$ as the input of our FA module. The second and the third ways preserve the information from TLM more effectively. The results are presented in Table LABEL:tab:tab6, indicating our method maintains a balance between prior knowledge and representation of information.
(a) Comparison between decoders
(b) Comparison between FA module
(c) the probability of being activated
(d) the $k$-th feature from mask decoder
Figure 7. (a) Comparison between one shared decoder and two separated decoders. (b) Comparison between lightweight and original feature aggregation module. (c) Experience on the probability of Pseudo Anomaly Generation being activated. (d) Ablation studies on the feature extracted from the $k$-th layer of mask decoder.
4.4.4. Ablation studies on the decoder and feature aggregation module
The image embeddings generated by the student encoder can closely approximate that of the original teacher encoder, leading to the conclusion that the utilization of distinct decoders for the TLM may not be imperative. as shown in Figure 7(a), We observe that the shared decoder, while economical in terms of parameters and computation time, achieves a satisfactory result with a slight decrease compared to employing two separated decoders. Furthermore, as indicated in Figure 7(b), it has been observed that reducing the depth of the FA module does not significantly impact the performance. As a result, an effort has been made to make this module more lightweight by reducing the channels from 256 to 128 and adjusting the dilation rate to [1, 1, 3], following a similar method in (Zhang et al., 2023d). These two lightweight steps mitigate memory costs associated with both training and inference, which is a factor of paramount importance in practical implementations.
4.4.5. Analysis of the probability of Pseudo Anomaly Generation
The probability of the Pseudo Anomaly Generation process being activated governs the extent to which the training datasets deviate from the original normal datasets. To be specific, a higher probability results in a larger proportion of abnormal images and a smaller proportion of normal images. The probability of 0.5 denotes an optimal balance between pseudo-anomalous images and normal images, as shown in Figure 7(c), achieving the best performance. Since the “good” samples in the test dataset, the probability of 1.0 lets normal images be “unseen” images during training, leading to a high false negative. We design experiments to verify this explanation in Section 4.4.9.
In Equation (3) and Equation (4), we define the layers of the feature extracted from the Mask Decoder as the $k$-th layer, which is also presented in Figure 2 (d): Element-wise Production. We conducted a series of experiments to investigate the impact of $k$, and the corresponding results are presented in Figure 7(d). We denoted the two layers as L1 and L2. We include image-level AUROC, pixel-level AUROC and PRO metrics to provide a comprehensive cooperation of the anomaly detection and localization capabilities of each layer.
@@ 4.4.7. Effect of the knowledge distillation framework
We investigate the importance of the fixed SAM teacher, i.e., only utilizing TLM during training and inference phases, as illustrated in Figure 8.In this scenario, we no longer use the knowledge distillation structure. Instead, we directly train TLM like a regular segmentation network, using the mask as a supervision signal to segment anomalous regions. The results, shown in Table 10, indicate that directly using a segmentation network to identify anomalous regions does not yield satisfactory results. In contrast, using a distillation model for differential learning tasks in TLM’s two students produces better outcomes.
4.4.8. Ablation studies on the distillation methods
SAM generates high-quality segmentation results for previously unseen images, making it suitable for logit distillation. As (Zhao et al., 2022) said, distillation methods often focus on distilling deep features from intermediate layers, while the significance of logit distillation is greatly overlooked. We compare the feature-based method with the logit-based method, which minimizes the focal loss and L1 loss between segmentation masks of teacher SAM and the student streams, as illustrated in Figure  9. We no longer use feature distillation; instead, we attempt to leverage SAM’s logit distillation capabilities. In this experiment, we directly use the segmentation results from SAM for distillation, aligning TLM’s segmentation results with the fixed SAM teacher. The experimental results, shown in Table 10, indicate that the performance of feature distillation is superior on average for our model, compared with directly utilizing segmentation results from a SAM-style architecture.
[…]
Method I-AUROC P-AUROC PRO
w/o SAM 89.40 90.02 79.19
Ours 99.05 98.26 94.92
[…]
Method I-AUROC P-AUROC PRO
Logit 93.45 91.65 80.38
Ours 99.05 98.26 94.92
@@ 4.4.8. Ablation studies on the distillation methods
In Table 12, we provide further quantitative results of the false negative rate (FNR) (Altman and Bland, 1994) to verify the analysis in the ablation study that the probability of 1.0 lets normal images be “unseen” images during training, leading to a high false negative. Furthermore, the TLM and FA are trained under the assumption that all pictures are abnormal, causing the model to ”think” all pictures are anomalies and incorrectly classify normal images as abnormal, thereby reducing performance. The false negative rate, i.e., miss rate, is defined to measure the proportion of instances that are incorrectly identified as negative when it is actually positive. The FNR is calculated as:
@@
Method Avg. Text.↓ Avg. Obj.↓ Avg.↓
@@
Parch 84.74 89.85 82.03
Scar 96.63 96.30 90.52
Union 97.49 97.16 91.06
@@
To evaluate the pseudo anomaly generation strategies, we present experimental results with the renowned strategies NSA (Schlüter et al., 2022) and Cutpaste (Li et al., 2021b) in Table 12. In most SOTA methods, i.e., DRAEM (Zavrtanik et al., 2021a), Prototypical (Zhang et al., 2023c) and DeSTSeg (Zhang et al., 2023b), the authors attempt to leverage synthetic anomalies generated by NSA, which are also utilized in our paper. Furthermore, based on NSA and Cutpaste,  (Kim et al., 2024; Yao et al., 2023b; Zhang et al., 2023c; Tien et al., 2023) employ a variety of complex or artificial operations that more closely resemble the real anomalies. Notably, SimpleNet (Liu et al., 2023) simulates anomalies by introducing Gaussian noise to normal features, a process that requires careful tuning of appropriate hyperparameters.
4.5. Discussion and Future Work
[…]
Additionally, Table 5 shows that our model excels in model parameters and inference flexibility. Large models with high computational demands are unsuitable for large-scale deployment, especially on robots or mobile devices, where slow inference speed also impacts production progress. Our method delivers superior results on average across four datasets while offering significant advantages in parameter efficiency and GPU memory usage, making it highly suitable for robotic and mobile device applications. Moreover, our STLM achieves the second-fastest inference speed and ranks third in GFLOPS. We attribute the discrepancy of STLM between fewer model parameters and longer inference time compared to DeSTSeg to the more mature software and hardware optimization for ResNet-like frameworks over Transformer-based architectures. It is a challenge that requires further study and could lead to performance improvements. In the future, we aim to optimize the inference speed of the image encoder and develop a model for multi-class anomaly detection using a simpler structure.
@@ 5. Conclusion
In industrial anomaly detection, the primary concerns in real-world applications include two aspects: model efficiency and mobile-friendliness. In this paper, we propose a novel framework called SAM-guided Two-stream Lightweight Model for unsupervised anomaly detection tasks, tailored to meet the demands of real-world industrial applications while capitalizing on the strong generalization ability of SAM. Our STLM effectively distills distinct knowledge from SAM into the Two-stream Lightweight Model (TLM), assigning different tasks to each stream. To be specific, one stream focuses on generating discriminative and generalized feature representations in both normal and anomalous locations, while the other stream reconstructs features without anomalies. Last, we employ a shared mask decoder and a feature aggregation module to generate anomaly maps. By the design of our Two-stream Lightweight Model, shared mask decoder and lightweight feature aggregation module, the calculation and the number of parameters of the networks can be significantly reduced and the inference speed can naturally be improved without noticeable accuracy loss. Experiments on the renowned four datasets demonstrate that our method achieves competitive results on anomaly detection tasks. Furthermore, it exhibits model efficiency and mobile-friendliness compared to state-of-the-art methods.